OpenStudy (anonymous):

Electric fields and electric potentials of spherical charge inside spherical conductor!

OpenStudy (anonymous):

So here is the approximate scheme of a figure and its E and V at different levels depending on R.

OpenStudy (anonymous):

R=5cm, ρ=2μC/m3 R1=24cm , thickness of a conducting sphere is 12cm. ( R2=36cm)

OpenStudy (anonymous):

E ( r > R2 ) = 9.42 / r2 (example) E (R1 < r < R2 ) =? E (R < r < R1 ) =? E (r < R ) =?

OpenStudy (anonymous):

V(r>R2) ─ V∞ =? \[V2= V(r>R2)| _{R _{b}}\] =? V(R < r < R1 ) ─ V(R1) =? V(r < R ) ─ V(R) =?

OpenStudy (anonymous):

inside the conductor there can be no charge so electric field is zero.all charges are distributed in the surface and so electric potentials are same all off the inside off the conductor and egual to surface potential

OpenStudy (anonymous):

what about the electric field inside the conducting sphere, there sure is a electric field. I'm stuck on this problem. it is a tricky one

OpenStudy (anonymous):

inside the conducting sphere is there andy empty space you can show your question exactly

OpenStudy (anonymous):

can you show your question exactly sorry

OpenStudy (anonymous):

it is a problem with few questions and it is in spanish. i attached the image of a figure . have you seen it?

OpenStudy (anonymous):

yes now ı have seen it and ı understand ı will doit in paper and ı will take a photo for you

OpenStudy (anonymous):

appreciate it! hope i will undesrtand conductors better with your solutions. :)

OpenStudy (anonymous):

ıf you cannot undesrtand anything you can ask

OpenStudy (anonymous):

thank you ! i am going trhough the notes now!

OpenStudy (anonymous):

looks legit !!! thank you for explanations too !!! wish i can give more than one medal

OpenStudy (anonymous):

you are welcome.I just had a final exam about these so ı will be a profession about these:)